OpenStudy (wolfe8):

Prove that (A × B)2 + (A · B)2 = A2B2 A and B are vectors.

OpenStudy (anonymous):

(AXB)^2 + (A.B)^2=A^2 B^2--->eq 1 But AXB=AB Sin theta (Magnitude) A.B=AB Cos theta put in eq 1 (AB Sin theta)^2 + (AB Cos theta)^2= A^2 B^2 A^2 B^2(Sin^2 theta+ Cos^2 theta)=A^2 B^2 ... But Sin^2 theta+Cos^2=1 A^2 B^2(1)=A^2 B^2 A^2 B^2=A^2 B^2 Hence proved.